# Regularized Langevin Dynamics for Combinatorial Optimization

### v2

@@ 2 Related Work
2.1 Simulated Annealing for CO
Simulated Annealing (SA) (Metropolis et al., 1953; Hastings, 1970; Neal, 1996; IBA, 2001) is a classic metaheuristic for combinatorial optimization (WANG et al., 2009; Bhattacharya et al., 2014; Tavakkoli-Moghaddam et al., 2007; Seçkiner & Kurt, 2007; Chen & Ke, 2004). However, traditional SA methods often require up to millions of sampling steps for convergence, which limits its applicability on large‐scale instances. The recent advances in discrete Markov Chain Monte Carlo (MCMC) offer the potential solutions by accelerating the convergence with locally-balanced proposals (Zanella, 2017). For example, GWG (Grathwohl et al., 2021) uses a first‐order approximation to estimate energy changes within the $1$‐Hamming‐ball neighborhood, steering each proposal toward high‐density regions. To mitigate local‐optima issues caused by small neighborhoods, PAS/PAFS (Sun et al., 2022) instead samples $d$ coordinates sequentially at each step. Their SA variant iSCO (Sun et al., 2023) achieves SOTA performance on various CO benchmarks. To address the inefficiency of sequential sampling in PAS/PAFS, Zhang et al. (2022) derive a discrete Langevin sampler from the continuous domain, assigning gradient‐informed change probabilities to all coordinates in parallel. However, direct transformation of the discrete Langevin sampler into an SA framework overlooks key domain differences, resulting in severe local‐optima issues. In this work, we build on these discrete MCMC developments by proposing a simple method that unifies parallel sampling with effective local‐optima escape.
@@ 3.1 Combinatorial Optimization Problem
where $a(\mathbf{x})$ stands for the target to optimize and $b(\mathbf{x})\geq 0$ corresponds to the amount of constraint violation (00 means no violation). In particular, we focus on the penalty form that can be written as
@@ 3.2 Langevin Dynamics
When $\mathbf{x}$ is a binary vector, we can obtain the flipping (changing the value of $\mathbf{x}_{i}$ from 00 to $1$, or $1$ to 00) probability:
@@ 4.2 Regularized Langevin Simulated Annealing
$s_{\tau}(\mathbf{x})=\log p_{\tau}(\mathbf{x})=-\frac{1}{\tau}\nabla H(\mathbf{x}).$ (12)
@@
This property allows us to efficiently regularize the SA algorithm with the $d$-th largest element in $\Delta$, denoted as $\Delta_{(d)}$. We then obtain the flipping probabilities by letting $\frac{1}{\alpha}=\frac{\Delta_{(d)}+\epsilon}{\tau}$, where $\epsilon\geq 0$ (e.g., $10^{-6}$) is optional:
$q(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})=\sigma\biggl{(}\frac{1}{2\tau}(\Delta_{i}-\Delta_{(d)}-\epsilon)\biggr{)}.$ (14)
@@
7:        $p\leftarrow\sigma\bigl{(}\frac{1}{2\tau}(\Delta_{i}-\Delta_{(d)}-\epsilon)\bigr{)}$
@@
It is worth noting that $k(\mathbf{x},\mathbf{x}^{\prime})=\exp\Bigl{(}-\frac{\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}^{2}}{2\alpha}\Bigr{)}$ is chosen here just for simplicity, but other strategies may perform just as well. For instance, one could omit the $-\frac{1}{\alpha}$ term and instead normalize the sigmoid outputs:
@@ 4.4 Connection to Normalized Gradient Descent
NGD is developed to address the vanishing/exploding gradient by normalizing the L2222 norm of the gradient for a scale-invariant update at each step. Our method, especially RLSA, could be treated as a discrete version of this approach by regularizing the Hamming distance between the solutions before and after the update. And both methods accelerate the descent under the smoothness condition.
@@ 5.1 Experimental Setup
  • RB and BA graphs: small (200200200200300300300300 nodes) and large (8008008008001200120012001200 nodes).
  • ER graphs: small (700700700700800800800800 nodes) and large (900090009000900011000110001100011000 nodes).
The large-scale ER graphs serve as a transfer‐testing set for models trained on small-scale ER. We append a suffix “-[ $n$$N$ ]” to each graph name to indicate its size range.
@@ Appendix B Comparison under Longer Runtime
iSCO (10×10$\times$10 ×) H $20.01$ 26.25m $40.47$ 1.87h $44.41$ 7.21m $378.56$ 11.03h
RLSA (10×10$\times$10 ×) H 20.10 6.98m 41.83 10.65m 45.05 2.92m 379.19 17.63m
@@
iSCO (10×10$\times$10 ×) H $18.97$ 8.81m $40.41$ 1.83h iSCO (10×10$\times$10 ×) H $734.62$ 1.20h $2960.23$ 43.98m
RLSA (10×10$\times$10 ×) H $18.97$ 3.14m 40.63 8.67m RLSA (10×10$\times$10 ×) H $734.62$ 4.07m 2968.59 10.25m
@@ 7 L = A@x
50 rr = torch.rand_like(x.data)

### v3

@@ 2 Related Work
2.1 Simulated Annealing Solvers for CO
Simulated Annealing (SA) (Metropolis et al., 1953; Hastings, 1970; Neal, 1996; IBA, 2001) is a classic metaheuristic for combinatorial optimization (WANG et al., 2009; Bhattacharya et al., 2014; Tavakkoli-Moghaddam et al., 2007; Seçkiner & Kurt, 2007; Chen & Ke, 2004). However, traditional SA methods often require up to millions of sampling steps for convergence, which limits its applicability on large‐scale instances. The recent advances in discrete Markov Chain Monte Carlo (MCMC) offer the potential solutions by accelerating the convergence with locally-balanced proposals (Zanella, 2017). For example, GWG (Grathwohl et al., 2021) uses a first‐order approximation to estimate energy changes within the 11‐Hamming‐ball neighborhood, steering each proposal toward high‐density regions. To mitigate local‐optima issues caused by small neighborhoods, PAS/PAFS (Sun et al., 2022) instead samples $d$ coordinates sequentially at each step. Their SA variant iSCO (Sun et al., 2023) achieves SOTA performance on various CO benchmarks. To address the inefficiency of sequential sampling in PAS/PAFS, Zhang et al. (2022) derive a discrete Langevin sampler from the continuous domain, assigning gradient‐informed change probabilities to all coordinates in parallel. However, direct transformation of the discrete Langevin sampler into an SA framework overlooks key domain differences, resulting in severe local‐optima issues. In this work, we build on these discrete MCMC developments by proposing a simple method that unifies parallel sampling with effective local‐optima escape.
@@ 3.1 Combinatorial Optimization Problem
where $a(\mathbf{x})$ stands for the target to optimize and $b(\mathbf{x})\geq 0$ corresponds to the amount of constraint violation (0 means no violation). In particular, we focus on the penalty form that can be written as
@@ 3.2 Langevin Dynamics
When $\mathbf{x}$ is a binary vector, we can obtain the flipping (changing the value of $\mathbf{x}_{i}$ from 0 to $1$, or $1$ to 0) probability:
@@ 4.2 Regularized Langevin Simulated Annealing
$s_{\tau}(\mathbf{x})=\nabla\log p_{\tau}(\mathbf{x})=-\frac{1}{\tau}\nabla H(\mathbf{x}).$ (12)
@@
This property allows us to efficiently regularize the SA algorithm with the $d$-th largest element in $\Delta$, denoted as $\Delta_{(d)}$. We then obtain the flipping probabilities by letting $\frac{1}{\alpha}=\frac{\Delta_{(d)}-\epsilon}{\tau}$, where $\epsilon\geq 0$ (e.g., $10^{-6}$) is optional:
$q(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})=\sigma\biggl(\frac{1}{2\tau}(\Delta_{i}-\Delta_{(d)}+\epsilon)\biggr).$ (14)
@@
7:   $p\leftarrow\sigma\bigl(\frac{1}{2\tau}(\Delta_{i}-\Delta_{(d)}+\epsilon)\bigr)$
@@
It is worth noting that $k(\mathbf{x},\mathbf{x}^{\prime})=\exp\Bigl(-\frac{\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}^{2}}{2\alpha}\Bigr)$ is chosen here just for simplicity, but other strategies may perform just as well. For instance, one could omit the $\frac{1}{\alpha}$ term and instead normalize the sigmoid outputs:
@@ 4.4 Connection to Normalized Gradient Descent
NGD is developed to address the vanishing/exploding gradient by normalizing the L22 norm of the gradient for a scale-invariant update at each step. Our method, especially RLSA, could be treated as a discrete version of this approach by regularizing the Hamming distance between the solutions before and after the update. And both methods accelerate the descent under the smoothness condition.
@@ 5.1 Experimental Setup
  • RB and BA graphs: small (200200300300 nodes) and large (80080012001200 nodes).
  • ER graphs: small (700700800800 nodes) and large (900090001100011000 nodes).
The large-scale ER graphs serve as a transfer‐testing set for models trained on small-scale ER. We append a suffix “-[ nnNN ]” to each graph name to indicate its size range.
@@ Appendix B Comparison under Longer Runtime
iSCO ($10\times$) H $20.01$ 26.25m $40.47$ 1.87h $44.41$ 7.21m $378.56$ 11.03h
RLSA ($10\times$) H 20.10 6.98m 41.83 10.65m 45.05 2.92m 379.19 17.63m
@@
iSCO ($10\times$) H $18.97$ 8.81m $40.41$ 1.83h iSCO ($10\times$) H $734.62$ 1.20h $2960.23$ 43.98m
RLSA ($10\times$) H $18.97$ 3.14m 40.63 8.67m RLSA ($10\times$) H $734.62$ 4.07m 2968.59 10.25m
@@ 7 L = A@x
50 $r$ = torch.rand_like(x.data)